OpenStudy (anonymous):

solve this by the addition method 3x+2y=-3 2x+3y=-2

OpenStudy (anonymous):

1) Multiply one of the equations by -1, therefore you get: 3x+2y=-3 -2x- 3y= 2 ----------- x - y = -1 2) Solve for x in terms of y x = y - 1 3) Substitute back into either equation 3(y-1) + 2y=-3 4) Solve for y y = 0 5) Substitute back into the other equation -2x-3(0) = 2 6) Solve for x x = -1

OpenStudy (anonymous):

Is that the addition method, don't know, I am asking?

OpenStudy (anonymous):

You need to multiply the first equation by 2 and the second equation by 3 to start. 6x + 4y = -6 6x + 9y = -6 ========= Subtract 0x + -5y = 0 dividing through results in y = 0 substitute that value of y in to either equation to get x = -1

OpenStudy (anonymous):

Sorry, thought you said subtraction method. Just multiply the second equation by -3 instead. 6x + 4y = -6 -6x - 9y = 6 ========== Add 0x - 5y = 0 etc.

OpenStudy (anonymous):

Hmm, a subtraction method...